OpenStudy (anonymous):

i feel so ashamed of asking this question lol but after several tries i give up ^^ [e^(x)*sqrt(e^(x))-e^(x)]/sqrt(e^(x))=e^(x)-sqrt(e^(x)) how??

OpenStudy (anonymous):

u just need to multiple both side with squr(e^X)

OpenStudy (anonymous):

note that squr (x)squr(x) =x and the same for any term

OpenStudy (experimentx):

\[ \frac{e^{x}*\sqrt{e^{x}}-e^{x}}{\sqrt{e^x}}=e^x-\sqrt{e^x}\] \[\sqrt{e^x}\sqrt{e^x} = e^x\]

OpenStudy (anonymous):

[e^(x)*sqrt(e^(x))-e^(x)]/sqrt(e^(x))=e^(x)-sqrt(e^(x)) * sqrt (e^x) e^(x)*sqrt(e^(x))-e^(x)]={e^(x)-sqrt(e^(x)} * sqrt (e^x)

OpenStudy (anonymous):

@0x58 are you comfortable with these explanation?

OpenStudy (anonymous):

yes but still can't get right right? cold you guys do it precisely??

OpenStudy (anonymous):

e^x * sqrt(e^x)-e^x = e^x ( sqrt e^x - 1)

OpenStudy (anonymous):

e^x ( sqrt e^x - 1) / sqrt e^x = sqrt e^x ( sqrt e^x -1) = 1 - sqrt e^x

OpenStudy (anonymous):

You seem get stuck at factor sqrt ( e^x), so I go the easier way!

OpenStudy (anonymous):

@0x58 Hope it's clear now!

OpenStudy (anonymous):

\[\sqrt{e ^{x}}\left( e ^{x}\sqrt{e ^{x}}-e ^{x} \right) \div \sqrt{e ^{x}} \sqrt{e ^{x}}=e ^{x}\sqrt{e ^{x}} \sqrt{e ^{x}}-e ^{x} \sqrt{e ^{x}} \div e ^{x}=e ^{x}\times e ^{x} - e ^{x} \sqrt{e ^{x}} \div e ^{x}\] is this right? if yes then what's next?

OpenStudy (anonymous):

I can't read like that! That's why I always separate them out: e^x ( sqrt e^x - 1) / sqrt e^x = sqrt e^x ( sqrt e^x -1) = 1 - sqrt e^x

OpenStudy (anonymous):

Oops = e^x - sqrt e^x ( typo at the last line)

OpenStudy (anonymous):

It's already extra difficult to read your post!

OpenStudy (anonymous):

what about this?? {sqrt(e ^(x))*[e^(x)*sqrt(e^(x))-e^(x)]} / sqrt(e ^(x))*sqrt(e ^(x)) ={e^(x)*sqrt(e^(x))*sqrt(e^(x))-e^(x)*sqrt(e^(x))} / e^(x) ={e^(x)*e^(x) - e^(x) * sqrt(e^(x))} / e^(x)

OpenStudy (anonymous):

@0x58 is it a new question ?

OpenStudy (anonymous):

no! just wanted to make sure that it's correct